### v1

# Under the hood of a carbon footprint calculator

Abstract.
We explain the mathematical theory of the Input-Output method for carbon footprints computations.
Introduction
Two main methods for calculating carbon footprints are the Input-Output (IO) method and the Lifecycle Analysis (LCA) method, also known as Craddle-to-Grave. Both aim to quantify the greenhouse gas (GHG) emissions of a product or service, but they differ in scope, data needs, and level of detail.
The Input-Output method is an economic approach developed in [L], that uses national or regional data on industries and their interdependencies. It calculates the carbon footprint by tracing the flow of goods and services across different sectors, estimating emissions based on economic input-output tables. It relies on aggregated data about economic sectors (e.g., agriculture, manufacturing) and their emissions, and estimates emissions associated with production, supply chains, and services by analyzing the economic connections between industries. This is the commonly accepted method for large-scale, economy-wide assessments is not data-intensive, relies on publicly available economic data. It will lack detail on specific products or processes and miss the emissions from smaller, niche activities.
The Lifecycle Analysis evaluates a product’s environmental impact over its entire lifecycle — from raw material extraction to disposal. It calculates emissions at each stage, considering both direct and indirect impacts. This provides a detailed, product-specific carbon footprint, identifies emissions hotspots and opportunities for reduction. It is however data-intensive and time-consuming, and the total complexity can lead to global inconsistencies, as the under or over counting of secondary emissions are only estimated.
The main difficulty in the concrete application of the Input-Output method is access to the relevant data: a recent Input-Output table (or transaction matrix) and a measure of environmental impact for each sector appearing in the table. The tables are quite available, see for Switzerland [IOCH]. The latter is harder to obtain, especially as it needs to be computed/collected in an homogeneous manner over the different sectors in order to make sense.
This note aims to explain the mathematical theory of the Input-Output method, and is based on the article by J. Kitzes [JK], the talk by J. Steinberger at the Climathics 2022 conference [Clim], and the third author’s talk for the Comission Romande de mathématiques in 2023.
1. A concrete question
To estimate the carbon footprint of a T-shirt for instance, one needs the footprint of the cotton all the way from production of the raw material up to the stitching of the T-shirt, all the washing during the lifespan of the T-shirt, and the disposal of the unusable T-shirt. But to estimate the cotton footprint only, one needs to estimate the transport of the water, the one of the pesticides used to protect the crop, the fertilizing of the field, all those requiring tractors and fuel, the fuel itself requiring transport and fuel, thus creating loops that are difficult to extricate. Moreover, once one estimates all the emissions, there’s no guarantee that all the emissions of all the objects that we so compute, will actually sum up to all the emissions that we actually produce. The Input-Output method takes the problem from the other end: it takes all the measured emissions, and then attributes it to different sectors, and then computes the carbon footprint of a goods from the textile industry according to its direct emissions, combined with all the indirect emissions as we shall now explain.
2. Basic Input-Output
This method models a closed economy with $n$ economical sectors, $S=(S_{1},\dots,S_{n})$, and comes in the following form:
$\begin{array}[]{c|cccc|cc}&S_{1}&S_{2}&\dots&S_{n}&D&T\\ \hline\cr S_{1}&c_{11}&c_{12}&\dots&c_{1n}&d_{1}&t_{1}\\ S_{2}&c_{21}&c_{22}&\dots&c_{2n}&d_{2}&t_{2}\\ \dots&\dots&\dots&\dots&\dots&\dots&\dots\\ S_{n}&c_{n1}&c_{n2}&\dots&c_{nn}&d_{n}&t_{n}\\ \hline\cr V&v_{1}&v_{2}&\dots&v_{n}&\\ T&t_{1}&t_{2}&\dots&t_{n}&\end{array}$
The $n\times n$ transaction matrix is $C=(c_{ij})\in M_{n}({\bf R}_{+})$ (positive coefficients only) has entry $c_{ij}$ the number representing the economical transactions from sector $S_{i}$ to $S_{j}$. The diagonal entries $c_{ii}$ then represent the internal gross revenue of the sector. The demand vector $D=(d_{1},\dots,d_{n})$ expresses the money spent by the population on each sector: $d_{i}$ is the amount bought from sector $S_{i}$, for $i=1,\dots,n$. The added value vector is given by $V=(v_{1},\dots,v_{n})$ and the total vector $T=(t_{1},\dots,t_{n})$ satisfy the following relations:
$t_{i}=\sum_{j=1}^{n}c_{ij}+d_{i}=v_{i}+\sum_{j=1}^{n}c_{ji}$
for $i=1,\dots,n$. The value $t_{i}$ represents the total output of the sector $S_{i}$, which is also the total input because the system is a closed circuit. Out of these vectors we define the technical coefficient matrix $A=(a_{ij})$ where
$a_{ij}=\frac{c_{ij}}{t_{j}}.$
For a vector $W=(w_{1},\dots,w_{n})$ we denote by $D_{W}$ the diagonal matrix with entries $w_{1},\dots,w_{n}$ in the diagonal, so that $A=CD_{T}^{-1}$. The scalar product of two vectors $X=(xi,\dots,x_{n})$ and $Y^{t}=(y_{1},\dots,y_{n})$ is given by
$X,Y\right>=\sum_{i=1}^{n}x_{i}y_{i}=XY$
where $X$ is a line vectors, namely $1\times n$ matrix, and $Y$ is a column vector, namely an $n\times 1$ matrix and $XY$ is the usual matrix multiplication so a real number.
3. Footprints in the Input-Output model
Economical activities have all sorts of footprints that are correlated to the income and the cost of things. To compute the footprint we need to have a emission vector $E=(e_{1},\dots,e_{n})$ where $e_{i}$ is the direct emission of sector $S_{i}$. It could be millions of tons of $CO_{2}$, measured or estimated directly above the industries of the sector, or square feet measured by the occupancy of the industries, or tons of plastic trash estimated by the collectivities, etc. Then $|E|=e_{1}+\cdots+e_{n}$ is the total footprint of the economical system. We compute the direct intensity of the print by $F=(f_{1},\dots,f_{n})$ with
$f_{i}=\frac{e_{i}}{t_{i}}$
each entry is the print per unit of money. Now, if we compute the inner product
$D,F\right>=\sum_{i=1}^{n}d_{i}f_{i}$
it is the footprint directly attributed to the consumers, but it doesn’t take into account all the intermediate exchanges that the sectors had to create the product. If you buy goods for $k$ dollars to industry $S_{i}$, then $kf_{i}$ indicates the direct footprint of the purchase, but not the total footprint of the goods. That one has to take into account the exchanges the sector $S_{i}$ had from the other ones. To do this, we need to compute the total intensity vector $X=(x_{1},\cdots,x_{n})$ that gives the total footprint of the sector $S_{i}$ by money amount of final consumer demand, given by
$X=F+FA+FA^{2}+\dots=F(I-A)^{-1}$
where $I$ denotes the $n\times n$ identity matrix (most matrices are invertible so let’s assume that one is).
Proposition 1.
The footprint is completely attributed to the consumers. Namely, with the above notations, $X,D\right>=|E|$.
Proof.
[…]
$X,D\right>=\left<F(I-A)^{-1},(I-A)T\right>=FT=\sum_{i=1}^{n}\frac{e_{i}}{t_{i}}t_{i}=|E|$
Remark 1.
Similarly one can define $Y=F(I-A^{t})^{-1}$ such that $Y,V\right>=|E|$ and that vector attributes the print to the added value of the different sector.
4. Limitations
Waste
In a closed economy, in order to take into account waste disposal, one needs a sector $S_{i}$ that runs the waste management. That sector has a large footprint since incinerators produce large quantities of $CO_{2}$ (of the order of 1kg $CO_{2}$ per kg of waste incinerated [Waste]), and is eventually attributed to the consumers through waste management taxes (in France the weight of waste is estimated at half a ton per person and per year). The model also only doesn’t discuss demand and waste, assuming that everything is bought and nothing gets wasted.
Added value
The computation doesn’t take into account the added value $V$, which is what the companies make, and the model attributes the footprint completely to the customers, through the vector $D$. Attributing the footprint to the added value created by the sectors would seem more natural to understand their responsibility, see Remark 1 above.
Inverses
Mathematically speaking, the main problem is that if $B$ is a matrix $\epsilon$-close to $A$ (say in terms of the coefficients), then the distance between $(I-A)^{-1}$ and $(I-B)^{-1}$ could be a priori be as big as we want, for instance if the matrices both have an eigenvalue close enough to 1. Since all the values are estimated, the model is probably not very stable in an optimization exercise.
Life Cycle Analysis (LCA)
To get a comprehensive and detailed view of the total carbon footprint of a given goods, one can examine each stage, including production, transportation, usage, and end-of-life processes. This method can estimate the footprint of goods with a detail that Input-Output can’t, but uses data-intensive estimates that need to be computed, with not total consistency guarantee whatsoever, and no way to estimate the inconsistency.